# Dusit Niyato, , Bo Ai

### v1

@@ V Numerical Results
In this section, we consider a CF mMIMO network, where all APs and UEs are randomly distributed in a $1\times 1\,\mathrm{km}^{2}$ area with the wrap-around scheme [5]. We assume that all AP-UE pairs have LoS paths and the large-scale fading pathloss coefficient between AP $m$ and UE $k$ is denoted by $\beta_{mk}$, which is modeled based on the COST 321 Walfish-Ikegami model as in [26, Eq. (24)]. The large-scale fading coefficient for the LoS and NLoS components between AP $m$ and UE $k$ can be denoted as $\beta_{mk}^{\mathrm{LoS}}=\frac{\kappa_{mk}}{\kappa_{mk}+1}\beta_{mk}$ and $\beta_{mk}^{\mathrm{NLoS}}=\frac{1}{\kappa_{mk}+1}\beta_{mk}$, respectively, where $\kappa_{mk}=10^{1.3-0.003d_{mk}}$ is the Rician $\kappa$ factor between AP $m$ and UE $k$ with $d_{mk}$ is the distance between AP $m$ and UE $k$ (taking the $11\,\mathrm{m}$ height difference into account). We consider that each AP is equipped with a uniform linear array (ULA) with $N$ antennas. The $n$-th element for $\bar{\mathbf{g}}_{mk}$ is $[\bar{\mathbf{g}}_{mk}]_{n}=\sqrt{\beta_{mk}^{\mathrm{LoS}}}e^{j2\pi(n-1)\sin\theta_{mk}{{\lambda}/{2}}}$, where $\lambda$ is the wavelength and $\theta_{mk}$ is the angle of arrival between UE $k$ and AP $m$. Note that the spatial correlation matrix $\check{\mathbf{R}}_{mk}$ is modeled based on the Gaussian local scattering model as in [5, Sec. 2.6]. Moreover, we have $\tau_{c}=200$, $\sigma^{2}=-94\,\mathrm{dBm}$, $p_{k}=200\,\mathrm{mW}$, and the bandwidth is $20\,\mathrm{MHz}$. The pilot allocation strategy follows from that of [26]. For each realization of the AP/UE locations, $1000$ channel realizations are considered.
@@
In Fig. 1, we investigate the average SE performance against the number of antennas per AP ($N$) for the fully centralized processing (FCP) scheme over different combining schemes. To validate the matching accuracy of the closed-form results for the FCP scheme with the Monte-Carlo simulation, we apply the UatF bound in (22) to evaluate the SE performance in this figure. As observed, with only channel statistics applied to design the C-OBE matrices, this scheme can achieve much better SE than the C-MR combining scheme. When $N=4$, the C-OBE combining can achieve $113.5\%$ SE improvement over the C-MR combining. More importantly, for $N=4$, there is only $33\%$ SE performance gap between the C-OBE and global optimal C-MMSE combining schemes. For the C-OBE and C-MR combining, markers $\circ$” generated by the closed-form results in Th. 1 and Th. 3 overlap with the curves generated by the Monte-Carlo simulation.
@@
Figure 1: Average SE measured by the UatF bound against the number of antennas per AP $N$ for the FCP scheme over different combining schemes with $M=20$, $K=20$, and $\tau_{p}=1$.
@@
Figure 3: Average SE performance measured by the standard bound against the number of UEs for the FCP scheme with $M=20$, $N=2$, and $\tau_{p}=\{1,5\}$.
Moreover, as examined, applying the standard capacity bound for the FCP scheme as [13, Proposition 1] to evaluate the SE performance with the C-OBE combining can achieve further SE improvement compared with the SE performance evaluated by the UatF bound in (5). Thus, in subsequent discussions, unless otherwise noted, to fully explore the potentials of the C-OBE scheme, we will utilize the standard capacity bound as [13, Proposition 1] to evaluate the SE performance for the FCP scheme. Fig. 2 shows the CDF curves of the SE per UE for the FCP scheme with different numbers of pilot signals. $\tau_{p}=1$ denotes that all $K$ UEs are assigned with the same pilot signal, where the pilot contamination is most severe. It can be found that the SE performance gap between the C-OBE and C-MMSE combining schemes for $\tau_{p}=1$ is much smaller than that of $\tau_{p}=5$. For instance, only $29\%$ average SE performance gap between the C-OBE and C-MMSE combining schemes for $\tau_{p}=1$, which is much smaller than $41.3\%$ for $\tau_{p}=5$. This observation demonstrates that the C-OBE scheme maintains robustness with severe pilot contamination, which is also clarified in Remark 7. However, when the pilot contamination is moderate, the C-MMSE combining can achieve significantly better performance than the C-OBE combining with the aid of high-quality channel estimates.
Figure 4: Average SE performance versus the number of APs over the EWDP and LSFD schemes with $K=20$, $N=4$, and $\tau_{p}=1$.
@@
Figure 8: Average SE performance for different OBE schemes over the Rician fading channels with or without phase-shifts with $M=20$, $K=20$, $N=4$, and $\tau_{p}=1$.
To further study the SE performance in the scenarios with pilot contamination, we investigate the average SE against the number of UEs in Fig. 3. As observed, the C-OBE scheme performs steadily in the scenario with severe pilot contamination. There is only $3.8\%$ performance gap between the SE with the C-OBE scheme with $\tau_{p}=1$ and $\tau_{p}=5$ for $K=30$. Moreover, the average SE performance gaps between the C-OBE and C-MMSE combining are $30.2\%$ and $48.4\%$ for $\tau_{p}=1$ and $\tau_{p}=5$ with $K=10$, respectively. These observations further demonstrate that the C-OBE combining is efficient to maintain robustness in scenarios with severe pilot contamination.
Next, we focus on the performance evaluation for the distributed processing schemes. [^3] For convenience, we let the terminology “equal weight distributed processing” and “LSFD” denote the performance analysis frameworks introduced in (26) and (16), respectively. Fig. 4 shows the average SE performance against $M$ over the equal weight distributed processing (EWDP) and LSFD schemes. The L-MMSE+LSFD scheme [^4] In the following, for convenience, we call “A+LSFD” as the processing scheme, where “A” combining scheme is applied at each AP and the optimal LSFD is implemented at the CPU. And “A” denotes that “A” combining scheme at each AP with no LSFD or just a simple equal weight scheme at the CPU. achieves the best performance among all schemes. However, the DG-OBE is also efficient since the performance gap between DG-OBE and L-MMSE+LSFD is only $7.5\%$ for $M=60$. Moreover, we find that the DG-OBE combining at each AP can achieve the same performance as the DG-OBE+LSFD scheme, which verifies the insights in Remark 5. This insightful observation demonstrates that the DG-OBE scheme can achieve an excellent SE performance with only the distributed combining applied at each AP and without the need to further implement the LSFD at the CPU.
@@
Next, we study the performance of the DL-OBE combining scheme. Fig. 6 shows the sum SE performance against $M$ under the EWDP and LSFD schemes. As observed, the DL-OBE scheme also performs well and there is only $9.1\%$ performance gap between the DL-OBE and L-MMSE+LSFD schemes for $N=4$. More importantly, it is interesting to find that the application of the optimal LSFD strategy at the CPU just brings marginal SE performance improvement than that of the pure DL-OBE scheme without further LSFD strategy, e.g. only $0.49\%$ SE improvement for the DL-OBE+LSFD scheme for $N=4$ compared with the DL-OBE scheme. Even though the DL-OBE scheme is locally designed based on local channel knowledge at each AP, it can achieve approximately the same performance as the DL-OBE+LSFD scheme, in which global statistical information is applied in the CPU to design the optimal LSFD strategy. This is because, due to the existence of random phase-shifts, many cross terms for different APs or different UEs become 0 and there is no further effective global information to be applied by the optimal LSFD strategy to enhance the performance. Besides, we notice that $\circ$” generated by the closed-form results in Th. 2 and Th. 5 match well with the curves generated by the Monte-Carlo simulation.
@@
As explained earlier in this paper, all the OBE matrices are designed based on the channel statistics. In CF mMIMO systems, the semi-deterministic LoS components are dominant, which are widely utilized for the design of OBE matrices. We will now showcase the important effects of the LoS components on the achievable SE performance for the OBE schemes. In Fig. 7, we show the sum SE performance against the “Rician range [^5] If the distance for a particular AP-UE link falls within the Rician range, this link is modelled using the Rician fading model. Conversely, if the distance exceeds this range, this link is modelled using the Rayleigh fading model without a LoS component. Thus, increasing the value of Rician range increases the likelihood that a particular link will include a LoS component.” for the L-MMSE+LSFD and DG-OBE schemes. We observe that the DG-OBE scheme performs more efficiently in the pure Rician scenario compared to the pure Rayleigh scenario. The performance gaps between the DG-OBE and L-MMSE+LSFD schemes are $13.5\%$ and $43.1\%$ over the Rician and Rayleigh scenarios, respectively. Moreover, as the Rician region increases, the achievable SE performance approaches to that of the pure Rician scenario. This is because continuously increasing Rician region can increase the number of Rician fading modelled links and thus enhances the system performance. Moreover, we observe that $\circ$” generated by the closed-form results in Th. 2 and Th. 4 match well with the curves generated by the Monte-Carlo simulation.
@@
To have a comprehensive overview of the proposed OBE schemes in this paper, we discuss the average SE performance for various centralized and distributed processing schemes over both the Rician channel model with or without phase-shifts as shown in Fig. 8. The bar diagrams and numerical values are generated by the Monte-Carlo simulation. We observe in the centralized processing case in Fig. 8 (a) that there is only $6.8\%$ average SE performance gap, measured by the standard capacity bound, between the C-OBE and the C-MMSE schemes in the scenario without phase-shifts, which is much smaller than the $28\%$ in the scenario with phase-shifts. Moreover, in the distributed processing case in Fig. 8 (b), we observe that the DG-OBE scheme can outperform the L-MMSE+LSFD scheme in the scenario without phase-shifts. Besides, in the scenario without phase-shifts, the DL-OBE+LSFD scheme can achieve $25.7\%$ average SE improvement compared to the DL-OBE scheme without the LSFD. This occurs because, in the scenario without phase-shifts, numerous global cross terms from different APs or UEs can be effectively utilized by the LSFD strategy at the CPU to further enhance performance. However, in the scenario with random phase-shifts, most of these cross terms become 00, significantly limiting the LSFD strategy’s potential to improve performance. Thus, the DL-OBE scheme can achieve approximately the same performance as the DG-OBE and DL-OBE+LSFD schemes in the scenario with phase-shifts.
@@
We will now extensively discuss all the OBE schemes proposed in this paper. Comparing Fig. 8 (a) and Fig. 8 (b), we find that, in the scenario with phase-shifts, the C-OBE scheme with the UatF bound achieves approximately the same performance as the distributed OBE schemes. As examined, in this scenario, the C-OBE matrix in (25) is a block diagonal one, where $M$ matrices with the $N\times N$ dimension each are located at the diagonal. These $M$ matrices are approximately equal to those of DG-OBE matrices from all $M$ APs. Thus, this block diagonal form of the C-OBE scheme makes the achievable SE for the FCP scheme, measured by the UatF bound in (5), approximately same as that of the DG-OBE scheme. In the scenario without phase-shifts, the performance gap between the C-OBE scheme under the UatF bound and the DG-OBE scheme is also marginal but the C-OBE scheme can achieve $6.5\%$ SE improvement over the DG-OBE scheme with the aid of the standard bound. Finally, by summarizing the above important observations, we provide some promising insights for all three OBE schemes:
@@
(1) C-OBE is advocated in the scenario with limited fronthaul: As demonstrated in [13, Fig. 6], the FCP scheme requires significantly less fronthaul signaling than the distributed processing scheme. The applications of both the DG-OBE and DL-OBE schemes necessitate extensive fronthaul signaling to transmit not only the soft data estimates but also channel statistics via fronthaul links. In contrast, the C-OBE scheme only requires the transmission of received pilot and data signals from all APs. Moreover, despite its higher computational complexity, the C-OBE scheme consistently achieves the best performance among all OBE schemes. Therefore, the C-OBE scheme is advocated in scenarios with limited fronthaul capacity, where all processing tasks can be centralized at the CPU.
@@
(3) DL-OBE is advocated in the scenario with phase-shifts in LoS components: In the scenarios with phase-shifts in LoS components, such as the mobile communication scenario and the scenario without perfect estimation and mitigation of phase-shifts, the DL-OBE scheme can approach the performance of the C-OBE and DG-OBE schemes with much lower computational complexity, thanks to the disappearance of most global cross terms caused by the random phase-shifts. Only local statistics information is required at each AP to design the DL-OBE matrix. Thus, the DL-OBE scheme is promising in the scenario with phase-shifts.

### v2

@@ V Numerical Results
In this section, we consider a CF mMIMO network, where all APs and UEs are randomly distributed in a $1\times 1\,\mathrm{km}^{2}$ area with the wrap-around scheme [5]. Unless otherwise stated, we assume that all AP-UE pairs have LoS paths and the large-scale fading pathloss coefficient between AP $m$ and UE $k$ is denoted by $\beta_{mk}$, which is modeled based on the COST 321 Walfish-Ikegami model in [29, Eq. (24)]. The large-scale fading coefficient for the LoS and NLoS components between AP $m$ and UE $k$ can be denoted as $\beta_{mk}^{\mathrm{LoS}}=\frac{\kappa_{mk}}{\kappa_{mk}+1}\beta_{mk}$ and $\beta_{mk}^{\mathrm{NLoS}}=\frac{1}{\kappa_{mk}+1}\beta_{mk}$, respectively, where $\kappa_{mk}=10^{1.3-0.003d_{mk}}$ is the Rician $\kappa$ factor between AP $m$ and UE $k$ with $d_{mk}$ is the distance between AP $m$ and UE $k$ (taking the $11\,\mathrm{m}$ height difference into account). We consider that each AP is equipped with a uniform linear array (ULA) with $N$ antennas. The $n$-th element for $\overline{\mathbf{g}}_{mk}$ is $[\overline{\mathbf{g}}_{mk}]_{n}=\sqrt{\beta_{mk}^{\mathrm{LoS}}}e^{j2\pi(n-1)\sin\theta_{mk}{{\lambda}/{2}}}$, where $\lambda$ is the wavelength and $\theta_{mk}$ is the angle of arrival between UE $k$ and AP $m$. Note that the spatial correlation matrix $\check{\mathbf{R}}_{mk}$ is modeled based on the Gaussian local scattering model in [5, Sec. 2.6]. Moreover, we have $\tau_{c}=200$, $\sigma^{2}=-94\,\mathrm{dBm}$, $p_{k}=200\,\mathrm{mW}$, and the bandwidth is $20\,\mathrm{MHz}$. The pilot allocation strategy follows from that of [29]. The sources of randomness embrace the random AP/UE locations and random fading realizations. More specifically, dozens or even hundreds of realizations of AP/UE locations are simulated and for each realization of the AP/UE locations, $1000$ channel realizations are considered. For the distributed processing scheme, the UatF capacity bound-based analysis framework is applied to evaluate the achievable SE performance. For the centralized processing scheme, here are two promising capacity bounds to evaluate the achievable SE performance: UatF capacity bound based on the channel statistics in (5) and the standard capacity bound based on the instantaneous channel state information in [16, Proposition 1]. We have mentioned which bound is applied in the following statements.
@@
In Fig. 1, we investigate the average SE per UE against the number of antennas per AP ($N$) for the fully centralized processing (FCP) scheme over different combining schemes. To validate the matching accuracy of the closed-form results for the FCP scheme with the Monte-Carlo simulation, we apply the UatF bound in (5) to evaluate the SE performance in this figure. As observed, with only channel statistics applied to design the C-OBE matrices, this scheme can achieve much better SE than the C-MR combining scheme. When $N=4$, the C-OBE combining can achieve $113.5\%$ SE improvement over the C-MR combining. More importantly, for $N=4$, there is only $33\%$ SE performance gap between the C-OBE and global optimal C-MMSE combining schemes. For the C-OBE and C-MR combining, markers $\circ$” generated by the closed-form results in Th. 1 and Th. 3 overlap with the curves generated by the Monte-Carlo simulation.
@@
Figure 1: Average SE per UE measured by the UatF bound against the number of antennas per AP for the FCP scheme over different combining schemes with $M=20$, $K=20$, and $\tau_{p}=1$.
Figure 2: CDF of the SE per UE measured by the standard, UatF, and genie-aided capacity bounds for the FCP scheme with $M=20$, $K=10$, $N=4$, and $\tau_{p}=1$.
@@
Figure 4: Average SE per UE measured by the standard bound against the number of UEs for the FCP scheme with $M=20$, $N=2$, and $\tau_{p}=\{1,5\}$.
Fig. 2 shows the cumulative distribution function (CDF) curves of the SE per UE, measured by the UatF capacity bound in (5), the standard capacity bound in [16, Proposition 1], and the genie-aided bound [^3] Note that the so-called genie-aided bound is the one that applies the perfect channel knowledge to compute the achievable SE, which is an upper value for the SE. in [36, Corollary 5.9] for the FCP scheme. We observe that when the C-MMSE combining scheme is applied, the performance gap between the SE performance evaluated by the UatF bound and that evaluated by the standard bound is negligible. This finding demonstrates that the UatF bound, which we utilized to design the C-OBE combining, is also efficient to depict the achievable SE performance. Moreover, there would be a further SE improvement by applying the standard capacity bound to evaluate the SE performance for the networks with the C-OBE combining scheme applied, compared with that of the UatF bound applied scenario. For instance, about $6.9\%$ average SE improvement can be achieved. Thus, in subsequent discussions, unless otherwise stated, to fully explore the potentials of the C-OBE scheme, we will utilize the standard capacity bound as [16, Proposition 1] to evaluate the SE performance for the FCP scheme. Meanwhile, there are about $8.1\%$ and $5.3\%$ average SE performance gaps between the SE performance evaluated by the standard bound and the genie-aided bound for the C-MMSE and C-OBE combining schemes, respectively. These gaps are acceptable since the genie-aided bound neglects the effect of both the bounding error and channel estimation error.
Fig. 3 shows the CDF curves of the SE per UE for the FCP scheme with different numbers of pilot signals. $\tau_{p}=1$ denotes that all $K$ UEs are assigned with the same pilot signal, where the pilot contamination is most severe. It can be found that the SE performance gap between the C-OBE and C-MMSE combining schemes for $\tau_{p}=1$ is much smaller than that of $\tau_{p}=5$. For instance, only $29\%$ average SE performance gap between the C-OBE and C-MMSE combining schemes for $\tau_{p}=1$, which is much smaller than $41.3\%$ for $\tau_{p}=5$. This observation demonstrates that the C-OBE scheme maintains robustness with severe pilot contamination, which is also clarified in Remark 11. However, when the pilot contamination is moderate, the C-MMSE combining can achieve significantly better performance than the C-OBE combining with the aid of high-quality channel estimates.
Figure 5: Average SE per UE versus the number of APs over the EWDP and LSFD schemes with $K=20$, $N=4$, and $\tau_{p}=1$.
@@
Figure 9: Average SE per UE for the distributed processing scheme over different channel estimators with $M=20$, $K=10$, $N=4$, and $\tau_{p}=1$.
@@
Figure 10: Average SE per UE for different OBE schemes over the Rician fading channels with or without phase-shifts with $M=20$, $K=20$, $N=4$, and $\tau_{p}=1$.
To further study the SE performance in the scenarios with pilot contamination, we investigate the average SE per UE against the number of UEs in Fig. 4. As observed, the C-OBE scheme performs steadily in the scenario with severe pilot contamination. There is only $3.8\%$ performance gap between the SE with the C-OBE scheme with $\tau_{p}=1$ and $\tau_{p}=5$ for $K=30$. Moreover, the average SE performance gaps between the C-OBE and C-MMSE combining are $30.2\%$ and $48.4\%$ for $\tau_{p}=1$ and $\tau_{p}=5$ with $K=10$, respectively. These observations further demonstrate that the C-OBE combining is efficient to maintain robustness in scenarios with severe pilot contamination.
Next, we focus on the performance evaluation for the distributed processing schemes. [^4] For convenience, we let the terminology “equal weight distributed processing” and “LSFD” denote the performance analysis frameworks introduced in (26) and (16), respectively. Fig. 5 shows the average SE per UE against $M$ over the equal weight distributed processing (EWDP) and LSFD schemes. The L-MMSE+LSFD scheme [^5] In the following, for convenience, we call “A+LSFD” as the processing scheme, where “A” combining scheme is applied at each AP and the optimal LSFD is implemented at the CPU. And “A” denotes that “A” combining scheme at each AP with no LSFD but just a simple equal weight scheme at the CPU. achieves the best performance among all schemes. However, the DG-OBE is also efficient since the performance gap between DG-OBE and L-MMSE+LSFD is only $7.5\%$ for $M=60$. Moreover, we find that the DG-OBE combining at each AP can achieve the same performance as the DG-OBE+LSFD scheme, which verifies the insights in Remark 7. This insightful observation demonstrates that the DG-OBE scheme can achieve an excellent SE performance with only the distributed combining applied at each AP and without the need to further implement the LSFD at the CPU.
@@
Next, we study the performance of the DL-OBE combining scheme. Fig. 7 shows the sum SE performance against $M$ under the EWDP and LSFD schemes. As observed, the DL-OBE scheme also performs well and there is only $9.1\%$ performance gap between the DL-OBE and L-MMSE+LSFD schemes for $N=4$. More importantly, it is interesting to find that the application of the optimal LSFD strategy at the CPU just brings marginal SE performance improvement than that of the pure DL-OBE scheme without further LSFD strategy, e.g. only $0.5\%$ SE improvement for the DL-OBE+LSFD scheme for $N=4$ compared with the DL-OBE scheme. Even though the DL-OBE scheme is locally designed based on local channel knowledge at each AP, it can achieve approximately the same performance as the DL-OBE+LSFD scheme, in which global statistical information is applied in the CPU to design the optimal LSFD strategy. This is because, due to the existence of random phase-shifts, many cross terms for different APs or different UEs become 0 and there is no further effective global information to be applied by the optimal LSFD strategy to enhance the performance. Besides, we notice that $\circ$” generated by the closed-form results in Th. 2 and Th. 5 match well with the curves generated by the Monte-Carlo simulation.
@@
As explained earlier in this paper, all the OBE matrices are designed based on the channel statistics. In CF mMIMO systems, the semi-deterministic LoS components are dominant, which are widely utilized for the design of OBE matrices. We will now show the important effects of the LoS components on the achievable SE performance for the OBE schemes. In Fig. 8, we show the sum SE performance against the “Rician range [^6] If the distance for a particular AP-UE link falls within the Rician range, this link is modelled using the Rician fading model. Conversely, if the distance exceeds this range, this link is modelled using the Rayleigh fading model without a LoS component. Thus, increasing the value of Rician range increases the likelihood that a particular link will include a LoS component.” for the L-MMSE+LSFD and DG-OBE schemes. We observe that the DG-OBE scheme performs more efficiently in the pure Rician scenario compared to the pure Rayleigh scenario. The performance gaps between the DG-OBE and L-MMSE+LSFD schemes are $13.5\%$ and $43.1\%$ over the Rician and Rayleigh scenarios, respectively. Moreover, as the Rician range increases, the achievable SE performance approaches to that of the pure Rician scenario. This is because continuously increasing Rician range can increase the number of Rician fading modelled links and thus enhances the system performance. We observe that $\circ$” generated by the closed-form results in Th. 2 and Th. 4 match well with the curves generated by the Monte-Carlo simulation.
Fig. 9 shows the average SE per UE for the distributed processing scheme when using different channel estimators. As observed, the distributed OBE schemes can also achieve excellent SE performance even without the prior knowledge of phase-shifts. Notably, the distributed OBE schemes can significantly outperform the L-MMSE combining-based scheme based on the LS estimator since the LS-based channel estimates show the poor ability to suppress interference. More interestingly, the distributed OBE schemes with the LS estimator can achieve similar SE performance compared to the distributed OBE schemes with the LMMSE estimator. This is because the distributed OBE schemes over both the LMMSE and LS estimators are linear to the $\mathbf{y}_{mk}^{p}$, and thus, can be uniformly constructed as $\mathbf{v}_{mk}=\mathbf{W}_{mk}\mathbf{y}_{mk}^{p}$. After optimizing the BE matrices $\mathbf{W}_{mk}$ based on different methodologies, the DG-OBE and DL-OBE schemes with the LS estimator are similar to those with the LMMSE estimator, which can achieve similar SE performance when using the LMMSE and LS estimators. Thus, under the scenario that the phase-shifts are not known or the scenario with the Rayleigh fading channel without the LoS component, if the OBE scheme is applied, only the low-complexity LS estimator is required to achieve similar SE performance as the computationally demanding MMSE-based channel estimators. Moreover, as observed, under the LMMSE and LS estimators, there is about $6.56\%$ SE improvement, which is much larger than that of the MMSE estimator, for the DL-OBE+LSFD scheme compared to the DL-OBE scheme since phase-shifts are not known at each AP, and thus, the LSFD strategy can slightly enhance the SE performance based on global statistics.
@@
To have a comprehensive overview of the proposed OBE schemes in this paper, we discuss the average SE per UE for various centralized and distributed processing schemes over both the Rician channel model with or without phase-shifts as shown in Fig. 10. The bar diagrams and numerical values are generated by the Monte-Carlo simulation. We observe in the centralized processing case in Fig. 10 (a) that there is only $6.8\%$ average SE performance gap, measured by the standard capacity bound, between the C-OBE and the C-MMSE schemes in the scenario without phase-shifts, which is much smaller than the $28\%$ in the scenario with phase-shifts. Moreover, in the distributed processing case in Fig. 10 (b), we observe that the DG-OBE scheme can outperform the L-MMSE+LSFD scheme in the scenario without phase-shifts. Besides, in the scenario without phase-shifts, the DL-OBE+LSFD scheme can achieve $25.7\%$ average SE improvement compared to the DL-OBE scheme without the LSFD. This occurs because, in the scenario without phase-shifts, numerous global cross terms from different APs or UEs can be effectively utilized by the LSFD strategy at the CPU to further enhance performance. However, in the scenario with random phase-shifts, most of these cross terms become 00, significantly limiting the LSFD strategy’s potential to improve performance. Thus, the DL-OBE scheme can achieve approximately the same performance as the DG-OBE and DL-OBE+LSFD schemes in the scenario with phase-shifts.
@@
We will now extensively discuss all the OBE schemes proposed in this paper. Comparing Fig. 10 (a) and Fig. 10 (b), we find that, in the scenario with phase-shifts, the C-OBE scheme with the UatF bound achieves approximately the same performance as the distributed OBE schemes. As examined, in this scenario, the C-OBE matrix in (25) is a block diagonal one, where $M$ matrices with the $N\times N$ dimension each are located at the diagonal. These $M$ matrices are approximately equal to those of DG-OBE matrices from all $M$ APs. Thus, this block diagonal form of the C-OBE matrix makes the achievable SE for the FCP scheme, measured by the UatF bound in (5), approximately same as that of the DG-OBE scheme. In the scenario without phase-shifts, the performance gap between the C-OBE scheme under the UatF bound and the DG-OBE scheme is also marginal but the C-OBE scheme can achieve $6.5\%$ SE improvement over the DG-OBE scheme with the aid of the standard bound. Finally, by summarizing the above important observations, we provide some promising insights for all three OBE schemes:
@@
(1) C-OBE is advocated in the scenario with limited fronthaul: As demonstrated in [16, Fig. 6], the FCP scheme requires significantly less fronthaul signaling than the distributed processing scheme. The applications of both the DG-OBE and DL-OBE schemes necessitate extensive fronthaul signaling to transmit not only the soft data estimates but also channel statistics via fronthaul links. In contrast, the C-OBE scheme only requires the transmission of received pilot and data signals from all APs. Moreover, despite its higher computational complexity, the C-OBE scheme consistently achieves the best performance among all OBE schemes. Therefore, the C-OBE scheme is advocated in scenarios with limited fronthaul capacity, where all processing tasks can be centralized at the CPU.
@@
(3) DL-OBE is advocated in the scenario with phase-shifts in LoS components: In the scenarios with phase-shifts in LoS components, such as the mobile communication scenario and the scenario without perfect estimation and mitigation of phase-shifts, the DL-OBE scheme can approach the performance of the C-OBE and DG-OBE schemes with much lower computational complexity, thanks to the disappearance of most global cross terms caused by the random phase-shifts. Only local statistics information is required at each AP to design the DL-OBE matrix. Thus, the DL-OBE scheme is promising in the scenario with phase-shifts in LoS components.